Grade 6Mechanics

please help find correct answer urgently

12:13 pmQ-the distance described during thelast half secondof upward motion of a body projected vertically upward is(help choose correct answer)

1-dependent upon the velocity of projection

2-always constant

3-dependent upon mass of body

4-dependent upon the time taken to reach maximum height

which is the one correct answer with reasoning

aryia
7 Years agoGrade 6

1 Answer

Arun
7 Years ago
say s is the distance covered by the object in upward direction.
s=ut-1/2  gt2 =ut-5t2
distance covered in (t-1) sec=s'=u(t-1)-5(t-1)2     
expand the prev identities.
so dist covered in last sec=s-s'=u-10t+5  ... (1)
 
Hence option B is correct
 
now,v=0=u-10t
so  u=10t  put this in equ 1:
so dist in last sec=10t-10t+5=5=constant